OpenStudy (anonymous):

8x^-1+5y^-1/-3x^-2_9y^-2= simplify complex fraction

OpenStudy (kc_kennylau):

What's the underline representing?

OpenStudy (kc_kennylau):

Did you miss out any parentheses?

OpenStudy (anonymous):

-3x^-2- 9y^-2

OpenStudy (anonymous):

Underline was a negative

OpenStudy (kc_kennylau):

Okay, but did you miss out any parentheses?

OpenStudy (anonymous):

No that is t math problem

OpenStudy (kc_kennylau):

That means is it \(\displaystyle\frac{8x^{-1}+5y^{-1}}{-3x^{-2}-9y^{-2}}\) or \(\displaystyle8x^{-1}+\frac{5y^{-1}}{-3x^{-2}}-9y^{-2}\)?

OpenStudy (anonymous):

The first one

OpenStudy (kc_kennylau):

That means you've missed out two pairs of parentheses.

OpenStudy (kc_kennylau):

Convert all the terms with negative indices into fraction first, i.e. 8x^(-1) => 8/x & etc

OpenStudy (kc_kennylau):

You should have \(\displaystyle\frac{\frac8x+\frac5y}{-\frac3{x^2}-\frac9{y^2}}\)

OpenStudy (anonymous):

Okay then multiply the whole thing by xtwo and ytwo

OpenStudy (kc_kennylau):

Yep you can do so, what do you get after doing so?

OpenStudy (kc_kennylau):

Draw it out or use equation editor, as your preference

OpenStudy (anonymous):

8xy^2+5x^2+y/3y^2-9x^2 drawing doesn't work well with my tablet but this is the solution I have on the answer key

OpenStudy (kc_kennylau):

Then you can just use plain text, but remember to put parentheses into appropriate position Oh God

OpenStudy (anonymous):

Thnks